OpenStudy (anonymous):

Mass of Cu needed to make 2.0 moles of Cu2O? May anyone show work?

OpenStudy (chmvijay):

can u tell me the atomic mass of Cu and molar mass of Cu2O

OpenStudy (chmvijay):

@xkittybunnyx ?/???

OpenStudy (anonymous):

atomic mass Cu = 63.5 and Molar mass Cu2O 143 thanks

OpenStudy (chmvijay):

so 143 gram of Cu2O has 2*63.5= 127 gram of Cu so 2 mole(2*143) of Cu2O will have =2*143*127 / 143 = find now

OpenStudy (rane):

mass = moles * molecular mass so your moles is 2 and molecualr mass which is known as (M) is = 63.546(Cu) + (16*2) of O2 = 2 * 95.456 = 190.912g

OpenStudy (anonymous):

Thanks but both of these answers or ways or different. May anyone else confirm?

OpenStudy (chmvijay):

u can consider mine one whee i converted whole thing from mole to weight